OpenStudy (anonymous):

Write the equation in slope-intercept form: 2x + 3y = 18 Answer y = 18 - 2x y = -2/3 x + 6 y = 3/2 x + 18 y = 18 - 3x

OpenStudy (anonymous):

just solve for y

OpenStudy (anonymous):

subtract 2x then divide by 3

OpenStudy (anonymous):

2x + 3y =18 3y= -2x + 18 y=-2/3x + 6

OpenStudy (anonymous):

medal please?

OpenStudy (anonymous):

there